### v1

@@ Abstract
Chain-of-Thought (CoT) [2, 3] guides large language models (LLMs) to reason step-by-step, and can motivate their logical reasoning ability. While effective for logical tasks, CoT is not conducive to creative problem-solving which often requires out-of-box thoughts and is crucial for innovation advancements. In this paper, we explore the Leap-of-Thought (LoT) abilities within LLMs — a non-sequential, creative paradigm involving strong associations and knowledge leaps. To this end, we study LLMs on the popular Oogiri game which needs participants to have good creativity and strong associative thinking for responding unexpectedly and humorously to the given image, text, or both, and thus is suitable for LoT study. Then to investigate LLMs’ LoT ability in the Oogiri game, we first build a multimodal and multilingual Oogiri-GO dataset which contains over 130,000 samples from the Oogiri game, and observe the insufficient LoT ability or failures of most existing LLMs on the Oogiri game. Accordingly, we introduce a creative Leap-of-Thought (CLoT) paradigm to improve LLM’s LoT ability. CLoT first formulates the Oogiri-GO dataset into LoT-oriented instruction tuning data to train pretrained LLM for achieving certain LoT humor generation and discrimination abilities. Then CLoT designs an explorative self-refinement that encourages the LLM to generate more creative LoT data via exploring parallels between seemingly unrelated concepts and selects high-quality data to train itself for self-refinement. CLoT not only excels in humor generation in the Oogiri game as shown in Fig. 1 but also boosts creative abilities in various tasks like “cloud guessing game” and “divergent association task”. These findings advance our understanding and offer a pathway to improve LLMs’ creative capacities for innovative applications across domains. The dataset, code, and models will be released online. https://github.com/sail-sg/CLoT.
Figure 1: Comaprison between (multimodal) large language model (LLM,   red) and its CLoT-integrated version (   blue) for Oogiri-style multimodal humor generation. According to the model input that can be image, text or both, there are three Oogiri tasks, “Image&Text to Text (IT2T)”, “Image to Text (I2T)”, and “Text to Text (T2T)”, where text can be English (EN), Chinese (CN), and Japanese (JP). “@” denotes translations in English. The baseline LLM is Qwen-VL [1]. While humor is subjective, these examples demonstrate CLoT’s leap-of-thought capacity of using excellent creative thinking to produce high-quality humor responses. See more examples in Appendix.
@@ 2 Related Works
(1) Oogiri game (大喜利) is a general term for a series of traditional Japanese comedy games. In ancient times, there were different types of Oogiri, such as actors performing sumo wrestling, telling ghost stories, etc. The modern Oogiri game mainly refers to one specific type known as Tonchi (頓智), typically presented in the format of game shows or intellectual quiz programs [28]. Players are provided with various multimodal contents, which can be simple questions, random images, etc., and are then prompted to come up with humorous, creative responses to achieve surprising comedic effects, as the examples are shown in Fig. 3. It is worth noting that the character “頓” in both Japanese and Chinese denote “sudden”, while “智” means “intelligence, insight or intuition”. This highlights the connection between the Oogiri game and the requirement for strong associative abilities in LoT, making Oogiri an ideal platform for exploring LoT capabilities within LLMs.
@@
(3) Computational humor is a branch of computational linguistics and artificial intelligence that uses computers in humor research [50], which encompasses various tasks, including humor detection [51, 52, 53, 54, 55, 56, 57, 58], humor interpretation [58, 59, 60, 61], and humor generation [62, 63, 64, 65, 66], etc. With the advancement of generative LLMs [4, 29, 1], humor generation has become a popular focus while humor generation still faces challenges such as insufficient punchlines [67] and limited in multimodal contexts [68, 69].
@@ 4.2 Explorative Self-Refinement
Qwen-VL+AIT (${}_{+\text{AIT (Ours)}}$ 7B 39.7 38.9 15.7 67.3 40.4+ ${}_{+\ 6.8}$ 38.8 30.5 15.7 62.3 36.8+ ${}_{+\ 6.7}$ 30.6 28.7 16.7 62.6 34.6+ ${}_{+\ 4.3}$
Qwen-VL+CLoT (${}_{+\text{CLoT (Ours)}}$ 7B 41.8 38.7 21.6 68.5 42.7+ ${}_{+\ 9.1}$ 39.8 35.1 22.7 64.4 40.5${}_{+10.4}$ 38.8 29.4 21.0 64.7 38.5+ ${}_{+\ 8.2}$
@@
Input: Input $\boldsymbol{I}$, CLoT-trained LLM $\mathcal{A}$, response number $n$ Output: Creative response $\boldsymbol{R}$. 1$\triangleright$ Creating the candidate responses 2:construct $n$ weakly-associated conditions $\{\boldsymbol{C}_{i}\}_{i=1}^{n}$ 3:$\{\boldsymbol{R}_{i}\}_{i=1}^{n}\leftarrow\mathcal{A}([\boldsymbol{I},\{\boldsymbol{C}_{i}\}_{i=1}^{n}])$ 4$\triangleright$ Choosing most creative response 5:Top-2 $\boldsymbol{R}^{\prime}_{1},\boldsymbol{R}^{\prime}_{2}\leftarrow\mathcal{A}([\boldsymbol{I},\{\boldsymbol{R}_{i}\}_{i=1}^{n}])$ with ranking ability 6:Best $\boldsymbol{R}\leftarrow\mathcal{A}([\boldsymbol{I},\boldsymbol{R}^{\prime}_{1},\boldsymbol{R}^{\prime}_{2}])$ with selection ability 7:return Best response $\boldsymbol{R}$.
@@ 4.3 CLoT Inference
CogVLM-17B+AIT (${}_{+\text{AIT (Ours)}}$ 7B 57.4 37.4 33.5 21.8 64.6 42.9${}_{+13.1}$ 55.4 46.5 26.4 18.2 64.4 42.2${}_{+10.3}$
CogVLM-17B+CLoT (${}_{+\text{CLoT (Ours)}}$ 7B 66.9 47.6 43.4 30.7 69.4 51.6${}_{+21.7}$ 64.8 52.9 33.6 21.8 68.6 48.3${}_{+16.4}$
@@ D.2 Machine Screening by LLM
After collecting raw Oogiri data as outlined in Appendix D.1, it is important to acknowledge that the Oogiri game, being a comedy game, may involve responses with biases or other offensive humor. Additionally, since the Oogiri game allows participation from any Internet user, the potential for encountering such issues grows with the game’s Internet dissemination. Therefore, effective filtering of the raw data becomes essential. Specifically, to prevent the inclusion of bias, violence, explicit content, offensive language, etc., we employ the multimodal language model Qwen-VL [1] as a checker for the initial screening of the raw data. This screening is performed by constructing safety-checking prompts. The design of the screening template for Qwen-VL is outlined as follows: Does the image or text contain content related to <Label>? Or the combination of image and text shows the metaphor related to <Label>? If so, kindly respond with “Yes”; otherwise, respond with “No.” Here is the text: <Text> where the tag <Label> represents the keyword (e.g., violence, explicit content, offensive language, etc.) and <Text> denotes the responses from per collected Oogiri sample. To further enhance the effectiveness of safety-checking, we additionally employ the <Label> utilized by NudeNet  [^3] https://github.com/notAI-tech/NudeNet, which includes a substantial number of keywords associated with Not Safe For Work (NSFW) content. Finally, after machine screening, the number of samples is reduced to about 160,000.
@@ E.2 The Details of Instruction Templates
Original Instruction Based on the image, think of a sentence that is unexpected and humorous. Let’s think outside the box. A satisfactory response is Image: <Image> <Response> Instruction with Condition Please carefully understand the image and give an answer that contains conditional words and is surprising and funny. Let’s think outside the box. A surprising and funny answer containing conditional word is Condition: <Condition> Image: <Image> <Response> Instruction for Ranking Please evaluate the degree of unexpected and humorous effect when each of the option contents is combined with the image. Options: A. <Content A>B. <Content B>C. <Content C>D. <Content D>E. <Content E>Response Format: Please respond in the format of ranking the humorousness of the options from high to low, for example, “1. A. xxx. 2. B. xxx. 3. C. xxx. 4. D. xxx. 5. E. xxx.”. Be sure to rank all five options. Let’s think outside the box. The result of ranking the options from most surprising and funny to least is Image: <Image> <Response> Instruction for 3T1 Selection Please select the option that, when combined with the image, creates an unexpected and humorous effect. Only one option meets the requirements. Options: A. <Content A>B. <Content B>C. <Content C>Response Format: Please respond in the format of “Option id. Option content”, for example, “A. xxx”. Let’s think outside the box. The satisfactory option is Image: <Image> <Response>
@@
Original Instruction Please carefully understand the provided question and come up with a surprising and humorous response. Question: <Question> Let’s think outside the box. A satisfactory response is <Response> Instruction with Condition Please carefully understand the question and give an answer that contains conditional words and is surprising and funny. Question: <Question> Let’s think outside the box. A surprising and funny answer containing conditional word is Condition: <Condition> <Response> Instruction for Ranking Please evaluate the degree of unexpected and humorous effect when each of the option contents is combined with the question. Question: <Question>Options: A. <Content A>B. <Content B>C. <Content C>D. <Content D>E. <Content E>Response Format: Please respond in the format of ranking the humorousness of the options from high to low, for example, “1. A. xxx. 2. B. xxx. 3. C. xxx. 4. D. xxx. 5. E. xxx.”. Be sure to rank all five options. Let’s think outside the box. The result of ranking the options from most surprising and funny to least is <Response> Instruction for 3T1 Selection Please select the option that, when combined with the question, creates an unexpected and humorous effect. Only one option meets the requirements. Question: <Question>Options: A. <Content A>B. <Content B>C. <Content C>
Response Format: Please respond in the format of “Option id. Option content”, for example, “A. xxx”. Let’s think outside the box. The satisfactory option is <Response>
where the tag <Question> denotes the text question of Oogiri data.
@@
Original Instruction In this image, there are sections of text that need to be completed, and the content to fill in is denoted by [MASK]. Let’s think outside the box and complete the [MASK] to make the response unexpectedly funny. A satisfactory response is Image: <Image> <Response> Instruction with Condition In this image, there are sections of text that need to be completed, and the content to fill in is denoted by [MASK]. Let’s think outside the box and complete the [MASK] with a response that contains conditional words and is surprising and funny. A surprising and funny response containing conditional word is Condition: <Condition> Image: <Image> <Response> Instruction for Ranking In this image, there are sections of text that need to be completed, and the content to fill in is denoted by [MASK]. Please evaluate the degree of unexpected and humorous effect when the options are the content of the [MASK]. Options: A. <Content A>B. <Content B>C. <Content C>D. <Content D>E. <Content E>Response Format: Please respond in the format of ranking the humorousness of the options from high to low, for example, “1. A. xxx. 2. B. xxx. 3. C. xxx. 4. D. xxx. 5. E. xxx.”. Be sure to rank all five options. Let’s think outside the box. The result of ranking the options from most surprising and funny to least is Image: <Image> <Response> Instruction for 3T1 Selection In this image, there are sections of text that need to be completed, and the content to fill in is denoted by [MASK]. Please select the option that, creates an unexpected and humorous effect when being the content of the [MASK]. Only one option meets the requirements. Options: A. <Content A>B. <Content B>C. <Content C>Response Format: Please respond in the format of “Option id. Option content”, for example, “A. xxx”. Let’s think outside the box. The satisfactory option is Image: <Image> <Response>
@@
From Table 1 in the main text, it is evident that the data volume for the IT2T category is significantly lower compared to the other two Oogiri game types. Consequently, there is a need to augment more data to enhance instruction tuning for IT2T. It is noteworthy that IT2T can be regarded as a form of masked language modeling (MLM) task [88, 89, 90]. Therefore, we contemplate achieving this objective by constructing MLM tasks for the I2T and T2T data types. The details are as follows: Mask Instruction for I2T Please carefully understand the provided image and complete the answer by replacing the [MASK] part to make the answer unexpectedly funny. Answer: <Answer with [MASK]>Let’s think outside the box. The content of [MASK] is Image: <Image> <Response> Mask Instruction for T2T Please carefully understand the provided question and complete the answer by replacing the [MASK] part to make the answer unexpectedly funny. Question: <Question>Answer: <Answer with [MASK]>Let’s think outside the box. The content of [MASK] is <Response> Here, we probabilistically replace nouns or verb phrases from the Oogiri answers with [MASK], and use the replaced Oogiri answer as <Answer with [MASK]>.
@@ E.4 The Extraction of Weakly-associated Conditions
For Japanese text, we use the specialized Japanese morphological analysis engine, Janome, to extract nouns from the Oogiri-GO corpus. The core implementation code is as follows, where tag is 名詞 (nouns).

### v3

@@ Abstract
Chain-of-Thought (CoT) [2, 3] guides large language models (LLMs) to reason step-by-step, and can motivate their logical reasoning ability. While effective for logical tasks, CoT is not conducive to creative problem-solving which often requires out-of-box thoughts and is crucial for innovation advancements. In this paper, we explore the Leap-of-Thought (LoT) abilities within LLMs — a non-sequential, creative paradigm involving strong associations and knowledge leaps. To this end, we study LLMs on the popular Oogiri game which needs participants to have good creativity and strong associative thinking for responding unexpectedly and humorously to the given image, text, or both, and thus is suitable for LoT study. Then to investigate LLMs’ LoT ability in the Oogiri game, we first build a multimodal and multilingual Oogiri-GO dataset which contains over 130,000 samples from the Oogiri game, and observe the insufficient LoT ability or failures of most existing LLMs on the Oogiri game. Accordingly, we introduce a creative Leap-of-Thought (CLoT) paradigm to improve LLM’s LoT ability. CLoT first formulates the Oogiri-GO dataset into LoT-oriented instruction tuning data to train pretrained LLM for achieving certain LoT humor generation and discrimination abilities. Then CLoT designs an explorative self-refinement that encourages the LLM to generate more creative LoT data via exploring parallels between seemingly unrelated concepts and selects high-quality data to train itself for self-refinement. CLoT not only excels in humor generation in the Oogiri game as shown in Fig. 1 but also boosts creative abilities in various tasks like “cloud guessing game” and “divergent association task”. These findings advance our understanding and offer a pathway to improve LLMs’ creative capacities for innovative applications across domains. The dataset, code, and models will be released online. https://zhongshsh.github.io/CLoT/.
Figure 1: Comparison between (multimodal) large language model (LLM,   red) and its CLoT-integrated version (   blue) for Oogiri-style multimodal humor generation. According to the model input that can be image, text or both, there are three Oogiri tasks, “Image&Text to Text (IT2T)”, “Image to Text (I2T)”, and “Text to Text (T2T)”, where text can be English (EN), Chinese (CN), and Japanese (JP). “@” denotes translations in English. The baseline LLM is Qwen-VL [1]. While humor is subjective, these examples demonstrate CLoT’s leap-of-thought capacity of using excellent creative thinking to produce high-quality humor responses. See more examples in Appendix.
@@ 2 Related Works
(1) Oogiri game {CJK*}UTF8goth(大喜利) is a general term for a series of traditional Japanese comedy games. In ancient times, there were different types of Oogiri, such as actors performing sumo wrestling, telling ghost stories, etc. The modern Oogiri game mainly refers to one specific type known as Tonchi {CJK*}UTF8goth(頓智), typically presented in the format of game shows or intellectual quiz programs [28]. Players are provided with various multimodal contents, which can be simple questions, random images, etc., and are then prompted to come up with humorous, creative responses to achieve surprising comedic effects, as the examples are shown in Fig. 3. It is worth noting that the character “{CJK*}UTF8goth頓” in both Japanese and Chinese denote “sudden”, while “{CJK*}UTF8goth智” means “intelligence, insight or intuition”. This highlights the connection between the Oogiri game and the requirement for strong associative abilities in LoT, making Oogiri an ideal platform for exploring LoT capabilities within LLMs.
@@
(3) Computational humor is a branch of computational linguistics and artificial intelligence that uses computers in humor research [50], which encompasses various tasks, including humor detection [51, 52, 53, 54, 55, 56, 57, 58], humor interpretation [58, 59, 60, 61], and humor generation [62, 63, 64, 65, 66, 67], etc. With the advancement of generative LLMs [4, 29, 1], humor generation has become a popular focus while humor generation still faces challenges such as insufficient punchlines [68] and limited in multimodal contexts [69, 70].
@@ 4.2 Explorative Self-Refinement
Qwen-VL+AIT (${}_{+\text{AIT (Ours)}}$ 7B 39.7 38.9 15.7 67.3 40.4+ 6.8 38.8 30.5 15.7 62.3 36.8+ 6.7 30.6 28.7 16.7 62.6 34.6+ 4.3
Qwen-VL+CLoT (${}_{+\text{CLoT (Ours)}}$ 7B 41.8 38.7 21.6 68.5 42.7+ 9.1 39.8 35.1 22.7 64.4 40.5+10.4 38.8 29.4 21.0 64.7 38.5+ 8.2
@@
Input: Input $\boldsymbol{I}$, CLoT-trained LLM $\mathcal{A}$, response number $n$
Output: Creative response $\boldsymbol{R}$.
1$\triangleright$ Creating the candidate responses
2:construct $n$ weakly-associated conditions $\{\boldsymbol{C}_{i}\}_{i=1}^{n}$
3:$\{\boldsymbol{R}_{i}\}_{i=1}^{n}\leftarrow\mathcal{A}([\boldsymbol{I},\{\boldsymbol{C}_{i}\}_{i=1}^{n}])$
4$\triangleright$ Choosing most creative response
5:Top-2 $\boldsymbol{R}^{\prime}_{1},\boldsymbol{R}^{\prime}_{2}\leftarrow\mathcal{A}([\boldsymbol{I},\{\boldsymbol{R}_{i}\}_{i=1}^{n}])$ with ranking ability
6:Best $\boldsymbol{R}\leftarrow\mathcal{A}([\boldsymbol{I},\boldsymbol{R}^{\prime}_{1},\boldsymbol{R}^{\prime}_{2}])$ with selection ability
7:return Best response $\boldsymbol{R}$.
@@ 4.3 CLoT Inference
CogVLM-17B+AIT (${}_{+\text{AIT (Ours)}}$ 7B 57.4 37.4 33.5 21.8 64.6 42.9+13.1 55.4 46.5 26.4 18.2 64.4 42.2+10.3
CogVLM-17B+CLoT (${}_{+\text{CLoT (Ours)}}$ 7B 66.9 47.6 43.4 30.7 69.4 51.6+21.7 64.8 52.9 33.6 21.8 68.6 48.3+16.4
@@ D.2 Machine Screening by LLM
After collecting raw Oogiri data as outlined in Appendix D.1, it is important to acknowledge that the Oogiri game, being a comedy game, may involve responses with biases or other offensive humor. Additionally, since the Oogiri game allows participation from any Internet user, the potential for encountering such issues grows with the game’s Internet dissemination. Therefore, effective filtering of the raw data becomes essential. Specifically, to prevent the inclusion of bias, violence, explicit content, offensive language, etc., we employ the multimodal language model Qwen-VL [1] as a checker for the initial screening of the raw data. This screening is performed by constructing safety-checking prompts. The design of the screening template for Qwen-VL is outlined as follows: {mdframed}[backgroundcolor=gray!8] Does the image or text contain content related to <Label>? Or the combination of image and text shows the metaphor related to <Label>? If so, kindly respond with “Yes”; otherwise, respond with “No.” Here is the text: <Text> where the tag <Label> represents the keyword (e.g., violence, explicit content, offensive language, etc.) and <Text> denotes the responses from per collected Oogiri sample. To further enhance the effectiveness of safety-checking, we additionally employ the <Label> utilized by NudeNet  [^3] https://github.com/notAI-tech/NudeNet, which includes a substantial number of keywords associated with Not Safe For Work (NSFW) content. Finally, after machine screening, the number of samples is reduced to about 160,000.
@@ E.2 The Details of Instruction Templates
{mdframed}
[backgroundcolor=gray!8] Original Instruction Based on the image, think of a sentence that is unexpected and humorous. Let’s think outside the box. A satisfactory response is Image: <Image> <Response> Instruction with Condition Please carefully understand the image and give an answer that contains conditional words and is surprising and funny. Let’s think outside the box. A surprising and funny answer containing conditional word is Condition: <Condition> Image: <Image> <Response> Instruction for Ranking Please evaluate the degree of unexpected and humorous effect when each of the option contents is combined with the image. Options:
A. <Content A>
B. <Content B>
C. <Content C>
D. <Content D>
E. <Content E>
Response Format: Please respond in the format of ranking the humorousness of the options from high to low, for example, “1. A. xxx. 2. B. xxx. 3. C. xxx. 4. D. xxx. 5. E. xxx.”. Be sure to rank all five options. Let’s think outside the box. The result of ranking the options from most surprising and funny to least is
Image: <Image> <Response> Instruction for 3T1 Selection Please select the option that, when combined with the image, creates an unexpected and humorous effect. Only one option meets the requirements. Options:
A. <Content A>
B. <Content B>
C. <Content C>
Response Format: Please respond in the format of “Option id. Option content”, for example, “A. xxx”. Let’s think outside the box. The satisfactory option is
Image: <Image> <Response>
@@
{mdframed}
[backgroundcolor=gray!8] Original Instruction Please carefully understand the provided question and come up with a surprising and humorous response. Question: <Question> Let’s think outside the box. A satisfactory response is <Response> Instruction with Condition Please carefully understand the question and give an answer that contains conditional words and is surprising and funny. Question: <Question> Let’s think outside the box. A surprising and funny answer containing conditional word is Condition: <Condition> <Response> Instruction for Ranking Please evaluate the degree of unexpected and humorous effect when each of the option contents is combined with the question. Question: <Question>
Options:
A. <Content A>
B. <Content B>
C. <Content C>
D. <Content D>
E. <Content E>
Response Format: Please respond in the format of ranking the humorousness of the options from high to low, for example, “1. A. xxx. 2. B. xxx. 3. C. xxx. 4. D. xxx. 5. E. xxx.”. Be sure to rank all five options. Let’s think outside the box. The result of ranking the options from most surprising and funny to least is
<Response> Instruction for 3T1 Selection Please select the option that, when combined with the question, creates an unexpected and humorous effect. Only one option meets the requirements. Question: <Question>
Options:
A. <Content A>
B. <Content B>
C. <Content C>
{mdframed}
[backgroundcolor=gray!8] Response Format: Please respond in the format of “Option id. Option content”, for example, “A. xxx”. Let’s think outside the box. The satisfactory option is <Response> where the tag <Question> denotes the text question of Oogiri data.
@@
[backgroundcolor=gray!8] Original Instruction In this image, there are sections of text that need to be completed, and the content to fill in is denoted by [MASK]. Let’s think outside the box and complete the [MASK] to make the response unexpectedly funny. A satisfactory response is Image: <Image> <Response> Instruction with Condition In this image, there are sections of text that need to be completed, and the content to fill in is denoted by [MASK]. Let’s think outside the box and complete the [MASK] with a response that contains conditional words and is surprising and funny. A surprising and funny response containing conditional word is Condition: <Condition> Image: <Image> <Response> Instruction for Ranking In this image, there are sections of text that need to be completed, and the content to fill in is denoted by [MASK]. Please evaluate the degree of unexpected and humorous effect when the options are the content of the [MASK]. Options:
A. <Content A>
B. <Content B>
C. <Content C>
D. <Content D>
E. <Content E>
Response Format: Please respond in the format of ranking the humorousness of the options from high to low, for example, “1. A. xxx. 2. B. xxx. 3. C. xxx. 4. D. xxx. 5. E. xxx.”. Be sure to rank all five options. Let’s think outside the box. The result of ranking the options from most surprising and funny to least is
Image: <Image> <Response> Instruction for 3T1 Selection In this image, there are sections of text that need to be completed, and the content to fill in is denoted by [MASK]. Please select the option that, creates an unexpected and humorous effect when being the content of the [MASK]. Only one option meets the requirements. Options:
A. <Content A>
B. <Content B>
C. <Content C>
Response Format: Please respond in the format of “Option id. Option content”, for example, “A. xxx”. Let’s think outside the box. The satisfactory option is
Image: <Image> <Response>
@@
From Table 1 in the main text, it is evident that the data volume for the IT2T category is significantly lower compared to the other two Oogiri game types. Consequently, there is a need to augment more data to enhance instruction tuning for IT2T. It is noteworthy that IT2T can be regarded as a form of masked language modeling (MLM) task [89, 90, 91]. Therefore, we contemplate achieving this objective by constructing MLM tasks for the I2T and T2T data types. The details are as follows: {mdframed}[backgroundcolor=gray!8] Mask Instruction for I2T Please carefully understand the provided image and complete the answer by replacing the [MASK] part to make the answer unexpectedly funny. Answer: <Answer with [MASK]>
Let’s think outside the box. The content of [MASK] is
Image: <Image> <Response> Mask Instruction for T2T Please carefully understand the provided question and complete the answer by replacing the [MASK] part to make the answer unexpectedly funny. Question: <Question>
Answer: <Answer with [MASK]>
Let’s think outside the box. The content of [MASK] is
<Response>
Here, we probabilistically replace nouns or verb phrases from the Oogiri answers with [MASK], and use the replaced Oogiri answer as <Answer with [MASK]>.
@@ E.4 The Extraction of Weakly-associated Conditions
For Japanese text, we use the specialized Japanese morphological analysis engine, Janome, to extract nouns from the Oogiri-GO corpus. The core implementation code is as follows, where tag is {CJK*}UTF8goth名詞 (nouns).